OpenStudy (destinykiara99):

Describe the transformation of the graph of f into the graph of g as either a horizontal or vertical stretch. f as a function of x is equal to the square root of x and g as a function of x is equal to 3 times the square root of x

OpenStudy (karatechopper):

So the difference between f(x)= sqrt(x) and g(x)=3sqrt(x) ?

OpenStudy (destinykiara99):

Yes

OpenStudy (karatechopper):

Well, in a horizontal stretch, lets devote a letter to where that stretch would occur in a function equation. Let's pick C. Horizontal stretch: y=f(Cx) Vertical stretch: y= Cf(x)

OpenStudy (karatechopper):

Make some sense?

OpenStudy (destinykiara99):

Yeah, that makes sense.

OpenStudy (karatechopper):

So tell me, by the sample format equations I just gave you, is the difference between the f(x) function and the g(x) function a horizontal or vertical stretch?

OpenStudy (destinykiara99):

I believe it's a horizontal stretch, right?

OpenStudy (karatechopper):

Take a look at where the 3 is.. Its outside the function, therefore vertical.

OpenStudy (destinykiara99):

Oh okay, yeah that makes sense. I'm sorry. I really don't understand this lesson so I appreciate you breaking it down.

OpenStudy (destinykiara99):

How would you describe the transformation? Would it be shifted up 3 units?

OpenStudy (destinykiara99):

@karatechopper

OpenStudy (karatechopper):

Sorry haah I got afk sometimes. It's basically like if you get a function like g(x)=3x Then you multiply the x constant by the 3

OpenStudy (karatechopper):

And then plot your point from there

OpenStudy (karatechopper):

We can stretch or compress it in the y-direction by multiplying the whole function by a constant.

OpenStudy (karatechopper):

If that makes more sense^

OpenStudy (destinykiara99):

Yes, thank you so much

OpenStudy (karatechopper):

https://www.mathsisfun.com/data/function-grapher.php?func1=sqrt(x)&func2=2*sqrt(1*(x+1))+1&xmin=-12&xmax=12&ymin=-8&ymax=8 Play around with this graph, it might provide a better visual! and np! Welcome to OpenStudy btw! :)